Three times the larger of two numbers is equal to four times the smaller. The sum of the numbers is 21. Find the numbers
Answer by bucky(2189)

The first sentence of the problem indicates that you are looking for two numbers. It also tells
you that there is a larger number (let's call it L) and as smaller number (let's call it S).
Finally, it tells you that 3 times the larger (3*L) is equal to 4 times the smaller (4*S).
In equation form this can be written as:
.
3*L = 4*S
.
The sum of the numbers is 21. Therefore, another equation is:
.
L + S = 21
.
Let's solve these two equations by substitution. We can start by solving the second equation
for L by subtracting S from both sides to leave just L by itself on the left side. When
you subtract S from both sides, the equation becomes:
.
L = 21 - S
.
Now let's substitute the right side of this equation for L in the first equation that we wrote.
The first equation is:
.
3*L = 4*S
.
substituting 21 - S for L changes the equation to:
.
3*(21 - S) = 4*S
.
Do the distributed multiplication on the left side by multiplying 3 times each of the terms
in the parentheses. 3 times 21 is 63 and 3 times -S is - 3*S. This changes the left side and
the equation becomes:
.
63 - 3*S = 4*S
.
Get rid of the -3*S on the left side by adding +3*S to both sides. On the left side, this
addition cancels the -3*S and on the right side the 3*S and 4*S add to give 7*S. So the
changed equation is:
.
63 = 7*S
.
Solve for S by dividing both sides of the equation by 7. This makes the answer for S:
.
S = 63/7 = 9
.
Now you know that the smaller number is 9. And since the total of the two numbers is
21, the larger number is 21 - 9 = 12.
.
Check: 3 times the larger number is 3*12 = 36. And 4 times the smaller number is 4*9 = 36.
This means that 3*L = 4*S
.
And L + S is 12 + 9 and that does equal 21, just as it should.
.
Therefore, the answers of 12 and 9 satisfy the problem and are, therefore, the correct answers.
.
Hope this helps you to understand the problem.
.